In a beauty salon, the waiting time to receive a haircut service is a random variable with a continuous uniform distribution. A calling for a new haircut service is every 10 minutes. (a) What is the probability that a customer waits for a new call more than 7 minutes? (b) What is the probability that a customer waits for a new call between 2 and 7 minutes?


1
Expert's answer
2022-03-23T15:44:09-0400

(a) "P(X>7)=\frac{10-7}{10}=0.3."


(b) "P(2<X<7)=\frac{7-2}{10}=0.5."